# Fuzzy Implicative Rules: A Unified Approach

### v1

@@ V Case Studio: Fuzzy implicative association rules
For data we have used five datasets from the UCI dataset repository [30]: abalone, iris, wdbc, magic and vehicle. Let us highlight that these datasets have both numeric and categorical variables. For numerical variables we have considered a fuzzy partition with three triangular fuzzy sets defined according the quantiles of the variable and for categorical variables crisp fuzzy partitions, according to the number of classes. For every database we have selected 0.3 as minimum fuzzy coverage and support and 0.8 as minimum fuzzy confidence. The results can be found in Table II.
@@
TABLE II: Results per dataset and pair of operators $(T,I)$. In each cell, it appears the number of fuzzy rules (top) and a tuple with the mean of the four quality measures FCoverage, FSupport, FConfidence and FWRAcc (bottom). Also, under the name of each dataset there is a summary of its specifications: (samples, features).
Dataset $(T,I)$ $(T_{\bm{P}},I_{Y})$ $(T_{\bm{LK}},I_{\bm{LK}})$ $(T_{\bm{P}},I_{\bm{(C_{\lambda}^{H})_{1,0}}})$ $(T_{\lambda}^{\bm{SS}},I^{\bm{SS}}_{k_{\lambda}})$
@@
(0.38,0.32,0.85,0.16)
13
(0.37,0.32,0.87,0.18)
vehicle
(846,19)
45
(0.36,0.31,0.87,0.18)
62
(0.36,0.33,0.9,0.19)
29
(0.37,0.31,0.85,0.18)
59
(0.37,0.33,0.9,0.19)
From Table II we can see that the choice of the pair $(T,I)$ affects the total number of rules and the quality metrics. Nonetheless, it is hard to get an idea of how different are the outputs without examining each set of output rules. Also, since all the measures introduced in Section III-C are defined in terms of the pair $(T,I)$, it is not adequate to compare any of these measures when different pairs $(T,I)$ are considered. To overcame this issue we propose a metric for comparing two sets of fuzzy rules using the Jaccard similarity index [31]. Specifically, if $SR_{1}=\{R_{1,1},\dots,R_{1,k}\}$ and $SR_{2}=\{R_{2,1},\dots,R_{2,k}\}$ are two sets of fuzzy rules, we define the percentage of similarity as $SR_{2}=\{R_{2,1},\dots,R_{2,k}\}$ are two sets of fuzzy rules, we define the percentage of similarity as
@@
where we have considered that $R_{1,l}=R_{2,\tilde{l}}$ if they involve the same target class and the same linguistic labels of the feature variables. In Figure 1 we can find the mean of similarity between the set of rules obtained in Table II for the different pair of operators.
Figure 1: Mean similarity between the set of rules obtained in Table II for different choices of $(T,I)$.
From the results in Figure 1 we can derive that the set of output rules obtained by different pairs of operators can be drastically different. Then, apart from modeling the support of a rules as a non-symmetric quantity, these results empirically show that fuzzy implicative rules capture different associations than conjunctive rules even if the same t-norm is used to model the conjunction.
@@ VI Conclusions
In this paper, we have revisited the concept of fuzzy implicative rules, and we have provided a solid theoretical framework for any algorithm interested in mining fuzzy rules modeled as logical conditionals. We have proved that for important characteristics like a monotone support or being the generalization of other frameworks it is necessary to introduce a new property which we have called (MTC). We have studied in depth this new property and we have provided different valid solutions. Although there are plenty of fuzzy implication functions available in the literature, we have exposed that finding meaningful solutions remains a challenging problem, especially if the user wants to model the evaluation of a rule as a non-symmetric quantity. Further, we have developed an open-source Python implementation of our framework for mining fuzzy implicative rules and we have shown that our approach provides distinct knowledge compared to others.
As future work, we want to further study the (MTC) property for obtaining more diverse fuzzy operators that can be considered to model fuzzy implicative rules. Also, we want to take into account fuzzy implicative rules in other rule mining techniques like subgroup discovery or exception rules. More generally, the role of fuzzy implication functions in the study of directional fuzzy rules could be further explored. Finally, we want to consider heuristics to improve the efficiency of mining fuzzy implicative rules.

### v2

@@ V-B Similarity between models
To assess whether fuzzy implicative rules yield knowledge that differs from other approaches, we conducted experiments on seven benchmark datasets from the UCI repository [31]: iris, wdbc, vehicle, abalone, magic, onlinenews and globalhousing. These datasets include both numerical and categorical variables. Unless specified otherwise, for the numerical variables, we constructed fuzzy partitions consisting of three triangular fuzzy sets defined according to the variable quantiles. For the categorical variables, we used crisp fuzzy partitions aligned with the number of categories. For every database, we have selected 0.3 as the minimum fuzzy coverage and support and 0.8 as the minimum fuzzy confidence, as we found it is a good balance between quality and number of rules (see Figure 3). As baselines for non-implicative fuzzy rule mining, we use the minimum $T_{\bm{M}}$ and the product $T_{\bm{P}}$ t-norms; for our proposed $(T,I)$ semantics we consider $(T_{\lambda}^{\bm{SS}},I^{\bm{SS}}_{k_{\lambda}})$ with $\lambda=-10$ and $(T_{\bm{LK}},I_{p})$ with $p=0.01$. The results can be found in Table III.
(a) iris
(b) wdbc
Figure 3: Effect of minimum fuzzy coverage on rule quality and set size (top 20% of rules by confidence). Left axis: mean confidence and mean support; right axis: number of rules. Fuzzy operators were fixed to $(I_{\bm{LK}},I_{p})$ with $p=0.01$.
@@
TABLE III: Results per dataset and model. In each cell, it appears the number of rules (top) and a tuple with the mean of the four quality measures Coverage, Support and Confidence. Also, under the name of each dataset there is a summary of its specifications: (samples, features).
Dataset Model $T_{\bm{P}}$ $T_{\bm{M}}$ $(T_{-10}^{\bm{SS}},I^{\bm{SS}}_{k_{-10}})$ $(T_{\bm{LK}},I_{0.01})$ Crisp
iris
(150,4)
17
(0.35,0.32,0.92)
19
(0.35,0.32,0.92)
15
(0.35,0.32,0.92)
32
(0.35,0.33,0.96)
39
(0.33,0.31,0.95)
wdbc
(569,11)
49
(0.36,0.32,0.9)
51
(0.36,0.34,0.93)
39
(0.37,0.32,0.88)
84
(0.36,0.34,0.93)
67
(0.37,0.34,0.93)
vehicle
(846,19)
101
(0.35,0.32,0.9)
118
(0.35,0.32,0.93)
90
(0.36,0.31,0.89)
262
(0.36,0.33,0.93)
294
(0.34,0.32,0.94)
@@
(0.38,0.32,0.85)
44
(0.37,0.34,0.91)
16
(0.36,0.32,0.88)
onlinenews
(39644,59)
6585322
(0.48,0.48,1)
6575127
(0.48,0.48,1)
6355817
(0.48,0.46,0.95)
6650640
(0.48,0.48,1)
320054
(0.42,0.43,1)
globalhousing
(200000,24)
5424
(0.39,0.39,1)
5382
(0.39,0.39,1)
5258
(0.40,0.38,0.96)
5847
(0.39,0.39,0.99)
549
(0.39,0.39,0.99)
Overall, the fuzzy models ($T_{\bm{P}}$, $T_{\bm{M}}$, $(T_{-10}^{\bm{SS}},I^{\bm{SS}}_{k_{-10}})$, and $(T_{\bm{LK}},I_{0.01})$) exhibit very similar behavior in terms of the quality measures Coverage, Support, and Confidence, which generally remain stable around 0.35, 0.32, and above 0.9, respectively. The main differences lie in the number of extracted rules: $(T_{\bm{LK}},I_{0.01})$ tends to produce the largest rule bases, particularly in smaller datasets such as iris and vehicle, while maintaining high confidence values. For larger datasets, such as onlinenews and globalhousing, all fuzzy models scale in a similar manner, generating thousands or even millions of rules with high confidence (close to 1) and moderate coverage and support. Regarding the crisp case, it generally produces a higher number of rules than the fuzzy approaches in small and medium-sized datasets, while preserving similar quality levels. However, for the largest datasets, the number of extracted crisp rules is considerably lower, suggesting a more restrictive or selective rule generation process.
Nonetheless, it is hard to get an idea of how different the outputs are without examining each set of output rules. Also, since all the measures introduced in Section III-C are defined in terms of the pair $(T,I)$, it is not adequate to compare any of these measures when different pairs $(T,I)$ are considered. To overcome this issue, we propose a metric for comparing two sets of fuzzy rules using the Jaccard similarity index [32]. Specifically, if $SR_{1}=\{R_{1,1},\dots,R_{1,k}\}$ and $SR_{2}=\{R_{2,1},\dots,R_{2,k}\}$ are two sets of fuzzy rules, we define the percentage of similarity as $SR_{2}=\{R_{2,1},\dots,R_{2,k}\}$ are two sets of fuzzy rules, we define the percentage of similarity as
@@
where we have considered that $R_{1,l}=R_{2,\tilde{l}}$ if they involve the same target class and the same linguistic labels of the feature variables. The resulting similarity measure can thus be applied to compare rule sets induced with different operators or fuzzy partitions, and it seamlessly accommodates the crisp case. Figure 4 reports the mean pairwise similarity among the rule sets obtained in Table III for the models under consideration.
Figure 4: Mean similarity between the set of rules obtained in Table III for different models.
In Figure 4 we can observe to main patterns. First, $T_{\bm{P}}$, $T_{\bm{M}}$, and $(T_{-10}^{\bm{SS}},I^{\bm{SS}}_{k_{-10}})$ show the largest mutual overlaps, despite $(T_{-10}^{\bm{SS}},I^{\bm{SS}}_{k_{-10}})$ being non-symmetric. Quantitatively, the off-diagonal similarities are high for $T_{\bm{P}}$$T_{\bm{M}}$ (0.83) and $T_{\bm{P}}$$(T_{-10}^{\bm{SS}},I^{\bm{SS}}_{k_{-10}})$ (0.8), and moderate for $T_{\bm{M}}$$(T_{-10}^{\bm{SS}},I^{\bm{SS}}_{k_{-10}})$ (0.7). This suggests that, under our thresholds, these three settings recover a largely overlapping core of rules—even when asymmetry is present. A plausible explanation is the structure of $(T_{-10}^{\bm{SS}},I^{\bm{SS}}_{k_{-10}})$, for which the support saturates to 1 on a specific region of the domain (see Eq. (7)). Second, $(T_{\bm{LK}},I_{p})$ stands apart: its similarity with the others is only moderate (0.58 with $T_{\bm{P}}$, 0.65 with $T_{\bm{M}}$) and lowest against $(T_{-10}^{\bm{SS}},I^{\bm{SS}}_{k_{-10}})$ (0.51), consistent with its explicitly implicative, directional semantics. Crisp behaves differently for a fundamental reason: the quality metrics (coverage, support, confidence) are computed on hard sets and do not account for fuzzy reasoning (no partial memberships or generalized modus ponens). Consequently, its rule set reflects a different inductive bias. This is visible in its low overlap with the fuzzy variants (0.35 with $T_{\bm{P}}$, 0.41 with $T_{\bm{M}}$, 0.28 with $(T_{-10}^{\bm{SS}},I^{\bm{SS}}_{k_{-10}})$, and 0.35 with $(T_{\bm{LK}},I_{p})$); its closest match is $T_{\bm{M}}$, which is more “crisp-like” in conjunction evaluation. Overall, Figure 4 indicates a shared rule core among $T_{\bm{P}}$, $T_{\bm{M}}$, and the (non-symmetric) $(T_{-10}^{\bm{SS}},I^{\bm{SS}}_{k_{-10}})$, while the implicative $(T_{\bm{LK}},I_{p})$ contributes distinct rule sets—retrieving information than other fuzzy models, as we discussed also in Section V-A.
@@ VI Conclusions
In this paper, we have revisited the concept of fuzzy implicative rules, and we have provided a solid theoretical framework for any algorithm interested in mining fuzzy rules modeled as logical conditionals. We have proved that for important characteristics like a monotone support or being the generalization of other frameworks, it is necessary to introduce a new property which we have called the monotonicity of the generalized modus ponens or (MTC). We have studied in depth this new property and we have provided different valid solutions. Although there are plenty of fuzzy implication functions available in the literature, we have shown that finding meaningful solutions remains a challenging problem, especially if the user wants to model the evaluation of a rule as a non-symmetric quantity. Also, we disclose that many pairs of operators lead to a symmetric expression of the generalized modus ponens, even if the fuzzy implication function involved is non-commutative. Further, although for some operators the support is non-symmetric, the set of rules obtained may not be much different from symmetric approaches. This shows the importance of the selection of the fuzzy operators depending on the patterns to be captured by the fuzzy rules.
Further, we have developed an open-source Python implementation of our framework for mining fuzzy implicative rules, proving that our approach yields distinct insights by introducing a novel similarity measure. As future work, we plan to further investigate the (MTC) property to design more diverse fuzzy operators for modeling fuzzy implicative rules. In addition, we aim to incorporate fuzzy implicative rules into other rule-mining techniques such as subgroup discovery and exception rules. More broadly, the role of fuzzy implication functions in directional fuzzy rules deserves deeper exploration, for example within fuzzy rule-based approaches like Mamdani or TSK systems. Finally, we intend to develop heuristics and more specific optimization techniques to enhance the efficiency of mining fuzzy implicative rules.